The class president prepared juice for the class. She prepared 5 liters of juice that was 3 parts pure juice and 7 parts water. This was too sweet. How many liters of water should she add so that the concentration is reduced to 25% and the sweetness is improved?
Answer by lwsshak3(11628)

let x=amt of water to add to get a 25% concentration
current concentration:
liters of juice = .3*5 =1.5 liters
liters of water = .7*5 =3.5 liters
new 25% concentration:
1.5/(1.5+3.5+x)=.25=1/4
5+x=6
x=1
ans:1 liter of water must be added to reduce the concentration of juice to 25%